A galvanic cell is constructed by coupling Ag and Cd electrodes dipped in 0.5M silver nitrate and 0.25M CdSO4 respectively at 250 C. Write the cell scheme cell reaction, and calculate emf of the cell. Given that SRPs of Ag and Cd are +0.80V and -0.40V respectively.


1
Expert's answer
2021-01-11T03:54:56-0500

the cell scheme:

А (–) Сd | Cd2+ || Ag| Ag (+) C

cell reaction:

(–) А: Сd – 2e- = Сd2+

(+) C: Ag+ 1e- = Ag


ECd/Cd2+ = –0.40 + (0.059*lg0.25)/2 = –0,418V

Ag/Ag+ = 0.80 + 0.059*lg0.5 = 0,782V

emf = 0,782 – ( –0,418) = 1,200V